# RoGUENeRF: A Robust Geometry-Consistent Universal Enhancer for NeRF

### v1

@@ Abstract
Recent advances in neural rendering have enabled highly photorealistic 3D scene reconstruction and novel view synthesis. Despite this progress, current state-of-the-art methods struggle to reconstruct high frequency detail, due to factors such as a low-frequency bias of radiance fields and inaccurate camera calibration. One approach to mitigate this issue is to enhance images post-rendering. 2D enhancers can be pre-trained to recover some detail but are agnostic to scene geometry and do not easily generalize to new distributions of image degradation. Conversely, existing 3D enhancers are able to transfer detail from nearby training images in a generalizable manner, but suffer from inaccurate camera calibration and can propagate errors from the geometry into rendered images. We propose a neural rendering enhancer, RoGUENeRF, which exploits the best of both paradigms. Our method is pre-trained to learn a general enhancer while also leveraging information from nearby training images via robust 3D alignment and geometry-aware fusion. Our approach restores high-frequency textures while maintaining geometric consistency and is also robust to inaccurate camera calibration. We show that RoGUENeRF substantially enhances the rendering quality of a wide range of neural rendering baselines, e.g. improving the PSNR of MipNeRF360 by 0.63dB and Nerfacto by 1.34dB on the real world 360v2 dataset.
@@ 1 Introduction
In this work, we present RoGUENeRF, a NeRF enhancer which is designed to improve the image quality of NeRF renderings while maintaining geometric consistency and being robust to inaccurate camera calibration. Firstly, we propose a novel combined 3D + 2D alignment and refinement mechanism which accurately finds correspondences between images from different camera viewpoints, even when one input is severely degraded, and can also compensate for inaccurate estimates of scene geometry and camera poses. Secondly, we propose a novel geometry-aware spatial attention module which regulates misaligned regions based on both camera distance and pixel-wise differences. Lastly, we propose a pre-training and fine-tuning strategy which learns a general geometry-consistent enhancement function that transfers well (i.e. fine-tunes as quickly as minute per scene) to different distributions of rendering degradations. We thoroughly evaluate our proposed approach on six different NeRF baselines across real world bounded and unbounded scenes from three datasets: LLFF[18], DTU [10] and 360v2 [3]. We show consistent improvements in PSNR, SSIM and LPIPS over every baseline and qualitatively demonstrate substantial improvements in image quality over baselines and state-of-the-art NeRF enhancers.
@@ 2.1 High Fidelity NeRF
Since the seminal work of Mildenhall et al. [19], there have been several works which aim to improve the fidelity of NeRF-based models. Some methods [11, 29] incorporate additional processing layers after the NeRF model, coupled with image quality specific loss functions. Roessle et al. [29] proposed GANeRF, using a 2D conditional generator trained adversarially to refine the rendered output. To maintain view-consistency, a discriminator is trained end-to-end together with the underlying NeRF model, requiring computationally expensive patch-based training. Combined with the time required to train the generator, this significantly increases optimization time per scene from 15 minutes for the underlying NeRF model to 58 hours. AlignNeRF [11] also incorporates additional processing layers trained end-to-end with the NeRF model by introducing a shallow convolutional network coupled with an alignment aware loss and relies on the shallowness of the enhancement network to maintain view consistency. Some methods tackle the problem of NeRF super-resolution [39, 8] which aim to reconstruct a high resolution (HR) image from a low resolution NeRF by using relevant patches from a HR reference frame. Other approaches attempt to improve the underlying differentiable rendering algorithm directly [2, 3, 37] to better tackle anti-aliasing effects, unbounded scenes or reflections and so are able to more accurately represent the visual characteristics of the scene.
Another popular line of work has been to address the slow training and inference time of NeRF by changing from an MLP to a faster representation, e.g. voxels [31], hash encodings [21, 40, 36], tensor decomposition [6], octrees [46], reducing training time per scene to as fast as a few seconds. ZipNeRF [4] marries the advantages of both high-fidelity and fast-training by combining voxel-grid data structures with cone-based rendering. Dynamic scenes captured as volumetric videos also pose a significant challenge for many existing NeRF-based models which otherwise perform well on static fixed scenes. Many methods model the additional time dimension [26, 14, 25, 15, 9] and are able to render short free viewpoint videos effectively but at the cost of lower image quality and temporal consistency. Despite huge progress in improving the fidelity of NeRFs, these models remain constrained by errors in training data and camera pose estimation leading to rendering artifacts in real world scenes.
@@ 4.2 Baseline Methods
To demonstrate the general applicability of our proposed enhancer, we extensively evaluate our method using six different baseline NeRF methods which together represent the evolution of neural rendering over the last few years. These include the current state-of-the-art with respect to image fidelity (MipNeRF360 [3], ZipNeRF [4]), training speed (Nerfacto [36], NeuS2++ an unbounded variant of NeuS2 [40]) and older seminal works (NeRF [19], TensoRF [6]). We also compare to state-of-the-art NeRF enhancers including Boosting View Synthesis [30], NeRFLiX [50] and NeRFLiX++ [49], an extension of NeRFLiX using a GAN-based degradation simulator. As there is no available code, we report results directly from the original works for [30] and [49]. For each dataset, we report results averaged across all scenes in Table 1, and provide a scene by scene breakdown of each method for the LLFF and 360v2 datasets in the supplementary.
Model Dataset PSNR (dB) SSIM LPIPS ZipNeRF 360v2 28.90 0.8367 0.1779 ZipNeRF + NeRFLiX 360v2 29.00 ( 0.10) 0.8317 ( 0.005) 0.2045 ( 15.0%) ZipNeRF + Ours 360v2 29.23 ( 0.33) 0.8465 ( 0.098) 0.1662 ( 6.6%) MipNeRF360 360v2 28.26 0.8050 0.2297 MipNeRF360 + NeRFLiX 360v2 28.44 ( 0.18) 0.8036 ( 0.001) 0.2441 ( 6.3%) MipNeRF360 + Ours 360v2 28.89 ( 0.63) 0.8302 ( 0.025) 0.1987 ( 13.5%) Nerfacto 360v2 26.11 0.7157 0.3266 Nerfacto + NeRFLiX 360v2 26.92( 0.81) 0.7410 ( 0.025) 0.3044 ( 6.8%) Nerfacto + Ours 360v2 27.45( 1.34) 0.7700 ( 0.054) 0.2670 ( 18.2%) NeuS2++ DTU 27.35 0.7587 0.4386 NeuS2++ + NeRFLiX DTU 27.40 ( 0.05) 0.7752 ( 0.017) 0.4167 ( 5.0%) NeuS2++ + Ours DTU 28.46 ( 1.11) 0.8237 ( 0.065) 0.3939 ( 10.2%) NeRF LLFF 26.57 0.8170 0.2389 NeRF + Boosting View Synthesis LLFF 27.08 ( 0.51) 0.8371 ( 0.020) - NeRF + NeRFLiX LLFF 27.17 ( 0.60) 0.8552 ( 0.038) 0.1695 ( 29.0%) NeRF + NeRFLiX++ LLFF 27.25 ( 0.68) 0.8580 ( 0.041) - NeRF + Ours LLFF 27.67 ( 1.10) 0.8713 ( 0.054) 0.1495 ( 37.4%) TensoRF LLFF 26.88 0.8432 0.1829 TensoRF + NeRFLiX LLFF 27.38 ( 0.50) 0.8652 ( 0.022) 0.1514 ( 17.2%) TensoRF + NeRFLiX++ LLFF 27.38 ( 0.50) 0.8660 ( 0.023) - TensoRF + Ours LLFF 27.58 ( 0.70) 0.8670 ( 0.024) 0.1494 ( 18.3%)
@@ 4.3 Quantitative and Qualitative Evaluation
We present quantitative results in Table 1. We show that our model achieves improvements in all metrics for all six NeRF baselines. On the 360v2 dataset, RoGUENeRF improves the PSNR of Nerfacto by 1.34dB, MipNeRF360 by 0.63dB and ZipNeRF by 0.33dB, and achieves corresponding reductions in LPIPS of 18.2%, 13.5% and 6.6% respectively. On the DTU and LLFF datasets, RoGUENeRF improves the PSNR of Neus2++ by 1.11dB, NeRF by 1.10d and TensoRF by 0.70dB, with corresponding reductions in LPIPS of 10.2%, 37.4% and 18.3% respectively. We consistently improve SSIM scores for all baselines and datasets. We also outperform state-of-the-art NeRF enhancers NeRFLiX and NeRFLiX++ in every setting. These results are reflected in the qualitative evaluation shown in Figures 1 and 3. We note that ZipNeRF already achieves very high fidelity in high-frequency regions, so the improvements from our enhancer are largely from denoising without over-smoothing. We show noticeable improvements over all other NeRF baseline models and NeRFLiX in restoring high-frequency details. We also present video results in the supplementary which demonstrate that RoGUENeRF has view-consistency in line with the baseline NeRF models, as opposed to NeRFLiX which is geometry-agnostic and suffers from flickering in high-frequency regions.
Model No Noise | Small Noise | Medium Noise | Large Noise Nerfacto 0.2956 0.4174 0.5249 0.5960 + Uformer 0.2914 0.4069 0.5139 0.6109 + NeRFLiX 0.2696 0.3629 0.4887 0.6090 + Ours 0.2077 0.3504 0.4772 0.5758
@@ 4.4 Robustness To Inaccurate Camera Calibration
We evaluate the robustness of our method to inaccurate camera calibration in Table 2. To simulate the effects of inaccurate camera pose estimation, we apply increasing levels of additive zero mean Gaussian noise to the camera poses estimated by COLMAP. We use different standard deviations for the rotation and position elements of the camera pose. The rotation standard deviation is expressed in degrees while position standard deviation is expressed as a unit of physical distance, with total scene size set to a bounding cube of length 2. For each noise level, we train Nerfacto using the noisy camera poses and evaluate the ability of our enhancer to effectively improve the perceptual quality of the noisy NeRF model. We outperform Uformer and NeRFLiX for every level of camera noise. For large noise, Uformer and NeRFLiX fail completely, while our approach still achieves an improvement in LPIPS and qualitatively produces a noticeable improvement over the noisy Nerfacto model, as shown in in Figure 4.
@@ 4.5 Ablation Study
We validate the importance of individual components of our method in Table 3. We conduct the study across all scenes in the LLFF dataset and present average results. Our combined contributions show an improvement of 1.1dB PSNR over NeRF and 0.93dB over a strong 2D baseline enhancer, Uformer [41] on which our proposed approach is based. We show a sharp improvement in LPIPS perceptual quality from our 3D alignment module (7.4%) , and a further sharp improvement on top of that from our non-rigid refinement (13.4%), which greatly improves the ability of our model to find accurate correspondences for high frequency textures and correct errors in geometry. We evaluate combinations of the following model components: UF: Baseline 2D enhancer Uformer which our approach is based on. NN: Nearest Neighbour Selection. 3D-A: Our 3D Alignment module comprising of feature reprojection using predicted depths and estimated camera poses. NRF: Our Non-Rigid Refinement module comprising of an iterative lightweight optical flow network which further aligns the reprojected features. GA: Our Geometric Attention module comprising a geometry-aware feature regulation mechanism. PT: The effect of pre-training our model. We conduct a more detailed ablation study on pre-training and fine-tuning in Table 4. Here we show two things; that our method can achieve large improvements in as little as 1 minute of fine-tuning; and that it is our complete combined contributions which allow our method to learn novel image degradations so quickly. Uformer+NN also has access to nearest neighbours from the training set, but as it is unaware of geometry, it is not capable of quickly leveraging the information from nearby viewpoints despite pre-training.
@@
4.6 Limitations and Societal Impact
Limitations A limitation of our method is the requirement to store all relevant training images. This could be prohibitive for very large scenes, especially at higher image resolutions. Secondly, although we are several times faster than the baseline NeRF methods, our approach does not yet achieve real time (i.e. 30fps) inference. Finally, the performance of our model is limited by the lack of availability of a large scale NeRF dataset for pre-training.
Societal Impact As NeRFs become more editable, our method could be misused to improve the photorealism of generated videos of people, places and objects. Currently our method requires ground truth images with a dense 3D coverage of the scene, so it is not trivial to use in the wild.
@@ 5 Conclusion
We have presented RoGUENeRF, a geometry-consistent NeRF enhancer which combines concepts from 3D and 2D vision to substantially improve the image quality of NeRF renderings in real world settings. Our model accurately finds correspondences between different camera views by performing 3D alignment and non-rigid refinement, while also being robust to errors in camera pose estimation and reducing reprojection artifacts with geometry-aware attention. RoGUENeRF achieves consistent improvements in PSNR, SSIM and LPIPS and qualitatively demonstrates significant improvements in image quality over six varied NeRF baselines and existing state-of-the-art NeRF enhancers across three challenging real world datasets. Our model demonstrates wide applicability and strong generalization, fine-tuning on a novel scene in as quickly as minute to learn the distribution of image degradations.
Supplementary Material
We present further detailed qualitative and quantitative evaluation of our method in Sections 6 and 10. We evaluate the amount of training data required to train our model compared to the state-of-the-art in Section 7 and compare inference speeds in Section 8. We present full implementation details of each component of our method in Section 9.
6 Video Results
We present video results of our proposed enhancer compared to NeRFLiX [50], MipNeRF360 [3], Nerfacto [36] and NeRF [19]. We demonstrate noticeable improvements over the baseline NeRF models and NeRFLiX, whilst retaining geometric consistency and restoring high-frequency textures. We encourage viewing the video results to fully appreciate the improvement in quality provided by our method.
7 Real World Data Efficiency
Real world NeRF datasets are typically small, sometimes only containing tens of images per scene. Exisiting NeRF enhancers like NeRFLiX [50] struggle to leverage such small datasets, requiring several orders of magnitude more training data. In contrast, our novel geometry-consistent alignment and fusion mechanisms allow our enhancer to be pre-trained and fine-tuned on relatively small datasets. We compare the results of training NeRFLiX on different dataset sizes with our method in Table 5. We show that NeRFLiX requires several orders of magnitude more simulated training data (Vimeo90K [44]) than our method to achieve any improvement at all over the baseline NeRF model, TensoRF [6]. When trained on only the real world LLFF dataset, NeRFLiX actually degrades the performance of TensoRF by 0.6dB. This is consistent with our findings that if we fine-tune NeRFLiX on a novel scene, quantitative performance gets worse. In contrast, our method can be pre-trained and fine-tuned on small datasets to achieve large improvements in performance.
Model Training Data # Training Frames PSNR TensoRF LLFF 262 26.88 TensoRF + NeRFLiX LLFF 262 26.28 ( 0.60) TensoRF + NeRFLiX LLFF + Vimeo90K-10% 45,490 26.71 ( 0.17) TensoRF + NeRFLiX LLFF + Vimeo90K-50% 226,404 27.08 ( 0.30) TensoRF + NeRFLiX LLFF + Vimeo90K-100% 452,546 27.38 ( 0.50) TensoRF + Ours LLFF 262 27.58 ( 0.70)
8 Inference Speed
We compare the inference speed of NeRFLiX and NeRFLiX++ [49] with our method for different image resolutions in Table 6. We measure inference speed on a single NVidia V100 GPU at full float32 precision using the Pytorch event time measurement function [23]. For NeRFLiX++, we report inference times directly from the original work. We note that for all quantitative and qualitative comparisons in this work, NeRFLiX performs test-time augmentation (TTA) by running 4 forward passes during inference with horizontally and vertically flipped inputs and averaging the result. We perform a single inference pass and do not use TTA. We show that RoGUENeRF is approximately 13 faster than NeRFLiX with TTA, 3 faster than NeRFLiX without TTA, and 3 slower than NeRFLiX++.
Model TTA Input Resolution Inference Time (ms) NeRFLiX Yes 512x512 4208 NeRFLiX No 512x512 1050 Ours No 512x512 337 NeRFLiX++ No 512x512 109 NeRFLiX Yes 1024x1024 18385 NeRFLiX No 1024x1024 4579 Ours No 1024x1024 1346 NeRFLiX++ No 1024x1024 433
9 Implementation Details
We present our entire pipeline in Figure 5. For each of the learnable components in our pipeline, we present detailed module architectures in Figures 6 and 7. We refer to a standard 2d convolutional layer as , and indicate the kernel size with , channel output size with , stride with and input/output dimensions for linear layers with /. We use zero padding and He initialization [7] for learnable weights. We use leaky relu with a slope of 0.01.
9.1 3D Alignment
We first extract features using the convolutional encoder block shown in Figure 6. We use the Kornia [28] geometry remap function with bilinear interpolation and zero padding to perform the 3D alignment of neighbouring training images to a novel test view. The inputs to the function are formulated as described in the main manuscript.
9.2 Iterative Refinement Network
We present the implementation details of our iterative refinement network in Figure 7. We perform our optical flow estimation at 8 downsampled resolution and apply bilinear upsampling to the estimated flow field to get back to full size. Our flow warping uses the PyTorch grid sample function with bilinear interpolation.
9.3 Geometry-Aware Attention
We present the architecture of our spatial attention and camera attention modules in Figure 6. We learn a bias for each convolutional and linear layer. Camera rotations are first converted to Euler angles before being input into the attention module.
9.4 Uformer
We use the official implementation of the Uformer [41]. We change the original architecture to accept a 64 channel input instead of 3 channels. Otherwise we keep the architecture identical to the original formulation.
10 Per-Scene Results
We present extended quantitative results in Tables 7 and 8 where we report results for each scene in the 360v2 [3] and LLFF [18] datasets. We show that RoGUENeRF consistently outperforms the baseline NeRF model and NeRFLiX on the majority of scenes across all metrics.
[…]

### v2

@@ Abstract
Recent advances in neural rendering have enabled highly photorealistic 3D scene reconstruction and novel view synthesis. Despite this progress, current state-of-the-art methods struggle to reconstruct high frequency detail, due to factors such as a low-frequency bias of radiance fields and inaccurate camera calibration. One approach to mitigate this issue is to enhance images post-rendering. 2D enhancers can be pre-trained to recover some detail but are agnostic to scene geometry and do not easily generalize to new distributions of image degradation. Conversely, existing 3D enhancers are able to transfer detail from nearby training images in a generalizable manner, but suffer from inaccurate camera calibration and can propagate errors from the geometry into rendered images. We propose a neural rendering enhancer, RoGUENeRF, which exploits the best of both paradigms. Our method is pre-trained to learn a general enhancer while also leveraging information from nearby training images via robust 3D alignment and geometry-aware fusion. Our approach restores high-frequency textures while maintaining geometric consistency and is also robust to inaccurate camera calibration. We show that RoGUENeRF substantially enhances the rendering quality of a wide range of neural rendering baselines, e.g. improving the PSNR of MipNeRF360 by 0.63dB and Nerfacto by 1.34dB on the real world 360v2 dataset. Project page: https://sib1.github.io/projects/roguenerf/
@@ 1 Introduction
In this work, we present RoGUENeRF, a NeRF enhancer which is designed to improve the image quality of NeRF renderings while maintaining geometric consistency and being robust to inaccurate camera calibration. Firstly, we propose a novel combined 3D + 2D alignment and refinement mechanism which accurately finds correspondences between images from different camera viewpoints, even when one input is severely degraded, and can also compensate for inaccurate estimates of scene geometry and camera poses. Secondly, we propose a novel geometry-aware spatial attention module which regulates misaligned regions based on both camera distance and pixel-wise differences. Lastly, we propose a pre-training and fine-tuning strategy which learns a general geometry-consistent enhancement function that transfers well (i.e. fine-tunes in under minutes per scene) to different distributions of rendering degradations. We thoroughly evaluate our proposed approach on six different NeRF baselines across real world bounded and unbounded scenes from three datasets: LLFF[22], DTU [11] and 360v2 [3]. We show consistent improvements in PSNR, SSIM and LPIPS over every baseline and qualitatively demonstrate substantial improvements in image quality over baselines and state-of-the-art NeRF enhancers.
@@ 2.1 High Fidelity NeRF
Since the seminal work of Mildenhall et al. [23], there have been several works which aim to improve the fidelity of NeRF-based models. Some methods [12, 34] incorporate additional processing layers after the NeRF model, coupled with image quality specific loss functions. Roessle et al. [34] proposed GANeRF, using a 2D conditional generator trained adversarially to refine the rendered output. To maintain view-consistency, a discriminator is trained end-to-end together with the underlying NeRF model, requiring computationally expensive patch-based training. Combined with the time required to train the generator, this significantly increases optimization time per scene from 15 minutes for the underlying NeRF model to 58 hours. AlignNeRF [12] also incorporates additional processing layers trained end-to-end with the NeRF model by introducing a shallow convolutional network coupled with an alignment aware loss and relies on the shallowness of the enhancement network to maintain view consistency. Some methods tackle the problem of NeRF super-resolution [49, 9] by using relevant patches from a HR reference frame. Other approaches attempt to improve the underlying differentiable rendering algorithm directly [2, 3, 47, 54] to better tackle anti-aliasing effects, unbounded scenes or reflections and can better model the characteristics of the scene. A further approach is to jointly optimize camera poses together with the NeRF to reduce geometric errors from incorrect poses [17, 5, 45, 20, 28].
Another popular line of work has been to address the slow training and inference time of NeRF by changing from an MLP to a faster representation, e.g. voxels [36], SDFs [46], MPIs [41, 42], hash encodings [26, 50, 44], tensor decomposition [7], octrees [56], reducing training time per scene to as fast as a few seconds. ZipNeRF [4] marries the advantages of both high-fidelity and fast-training by combining voxel-grid data structures with cone-based rendering. Dynamic scenes captured as volumetric videos also pose a significant challenge for many existing neural rendering models which otherwise perform well on static fixed scenes. Many methods model the additional time dimension [32, 16, 31, 18, 10, 39, 25, 8] and are able to render short free viewpoint videos effectively but at the cost of lower image quality and temporal consistency. Despite huge progress in improving the fidelity of NeRFs, these models remain constrained by errors in training data and camera pose estimation leading to rendering artifacts in real world scenes.
@@ 4.2 Baseline Methods
To demonstrate the general applicability of our proposed enhancer, we extensively evaluate our method using six different baseline NeRF methods which together represent the evolution of neural rendering over the last few years. These include the current state-of-the-art with respect to image fidelity (MipNeRF360 [3], ZipNeRF [4]), training speed (Nerfacto [44], NeuS2++ an unbounded variant of NeuS2 [50]) and older seminal works (NeRF [23], TensoRF [7]). We compare to state-of-the-art NeRF enhancers including Boosting View Synthesis [35], NeRFLiX [60] and NeRFLiX++ [59], an extension of NeRFLiX using a GAN-based degradation simulator. For each dataset, we report results averaged across all scenes in Table 1. We provide per-scene results of each method and also compare to AligNeRF [12] on outdoor scenes in the supplementary. As there is no available code, we report results directly from the original works for [12], [35] and [59].
Model Dataset PSNR (dB) SSIM LPIPS ZipNeRF 360v2 28.90 0.8367 0.1779 ZipNeRF + NeRFLiX 360v2 29.00 ( 0.10) 0.8317 ( 0.005) 0.2045 ( 15.0%) ZipNeRF + Ours 360v2 29.23 ( 0.33) 0.8465 ( 0.098) 0.1662 ( 6.6%) MipNeRF360 360v2 28.26 0.8050 0.2297 MipNeRF360 + NeRFLiX 360v2 28.44 ( 0.18) 0.8036 ( 0.001) 0.2441 ( 6.3%) MipNeRF360 + Ours 360v2 28.89 ( 0.63) 0.8302 ( 0.025) 0.1987 ( 13.5%) Nerfacto 360v2 26.11 0.7157 0.3266 Nerfacto + NeRFLiX 360v2 26.92( 0.81) 0.7410 ( 0.025) 0.3044 ( 6.8%) Nerfacto + Ours 360v2 27.45( 1.34) 0.7700 ( 0.054) 0.2670 ( 18.2%) NeuS2++ DTU 27.35 0.7587 0.4386 NeuS2++ + NeRFLiX DTU 27.40 ( 0.05) 0.7752 ( 0.017) 0.4167 ( 5.0%) NeuS2++ + Ours DTU 28.46 ( 1.11) 0.8237 ( 0.065) 0.3939 ( 10.2%) NeRF LLFF 26.57 0.8170 0.2389 NeRF + Boosting View Synthesis† LLFF 27.08 ( 0.51) 0.8371 ( 0.020) - NeRF + NeRFLiX LLFF 27.17 ( 0.60) 0.8552 ( 0.038) 0.1695 ( 29.0%) NeRF + NeRFLiX++† LLFF 27.25 ( 0.68) 0.8580 ( 0.041) - NeRF + Ours LLFF 27.67 ( 1.10) 0.8713 ( 0.054) 0.1495 ( 37.4%) TensoRF LLFF 26.88 0.8432 0.1829 TensoRF + NeRFLiX LLFF 27.38 ( 0.50) 0.8652 ( 0.022) 0.1514 ( 17.2%) TensoRF + NeRFLiX++† LLFF 27.38 ( 0.50) 0.8660 ( 0.023) - TensoRF + Ours LLFF 27.58 ( 0.70) 0.8670 ( 0.024) 0.1494 ( 18.3%)
@@ 4.3 Quantitative and Qualitative Evaluation
We present quantitative results in Table 1. We show that our model achieves improvements in all metrics for all six NeRF baselines. On the 360v2 dataset, RoGUENeRF improves the PSNR of Nerfacto by 1.34dB, MipNeRF360 by 0.63dB and ZipNeRF by 0.33dB, and achieves corresponding reductions in LPIPS of 18.2%, 13.5% and 6.6% respectively. On the DTU and LLFF datasets, RoGUENeRF improves the PSNR of Neus2++ by 1.11dB, NeRF by 1.10dB and TensoRF by 0.70dB, with corresponding reductions in LPIPS of 10.2%, 37.4% and 18.3% respectively. We consistently improve SSIM scores for all baselines and datasets. We also outperform state-of-the-art NeRF enhancers NeRFLiX and NeRFLiX++ in every setting. These results are reflected in the qualitative evaluation shown in Figures 1 and 3. We note that ZipNeRF already achieves very high fidelity in high-frequency regions, so the improvements from our enhancer are largely from denoising without over-smoothing. We show noticeable improvements over all other NeRF baseline models and NeRFLiX in restoring high-frequency details. We also present video results in the supplementary which demonstrate that RoGUENeRF has view-consistency in line with the baseline NeRF models, as opposed to NeRFLiX which is geometry-agnostic and suffers from flickering in high-frequency regions.
Model No Noise | Small Noise | Medium Noise | Large Noise Nerfacto 66.50 53.33 41.65 32.43 Nerfacto + Ours 72.32 (5.8) 62.59 (9.3) 58.94 (17.3) 51.60 (19.2) ZipNeRF 71.38 63.29 58.67 49.50 ZipNeRF + Ours 72.06 (0.7) 69.76 (6.5) 65.03 (6.4) 52.39(2.9)
@@ 4.4 Robustness To Inaccurate Camera Calibration
We evaluate the robustness of our method to inaccurate camera calibration in Table 2. To simulate the effects of inaccurate camera pose estimation, we apply increasing levels of additive zero mean Gaussian noise to the camera poses estimated by COLMAP. The rotation standard deviation is expressed in degrees while position standard deviation is expressed as a unit of physical distance, with total scene size set to a bounding cube of length 2. For each noise level, we train Nerfacto and ZipNeRF using the noisy camera poses and evaluate the ability of our enhancer to improve the perceptual quality of the noisy NeRF baselines. Training with incorrect camera poses introduces pixel shifts between rendered images and ground truth, so we assess performance with MUSIQ [14], a SOTA no-reference image quality assessment metric. We show that our improvements over Nerfacto and ZipNeRF are larger when using noisy camera poses compared to the no noise setting. This is reflected in Figure 4 where we qualitatively evaluate the medium noise setting. The baselines suffer a large drop in image quality while RoGUENeRF is robust to inaccurate poses and is able to achieve noticeable improvements over the noisy baselines.
@@ 4.5 Ablation Study
We validate the importance of individual components of our method in Table 3. We conduct the study across all scenes in the LLFF dataset and present average results. Our combined contributions show an improvement of 1.1dB PSNR over NeRF and 0.93dB over a strong 2D baseline enhancer, Uformer [51] on which our proposed approach is based. We show a sharp improvement in LPIPS perceptual quality from our 3D alignment module (7.4%), and a further improvement from our non-rigid refinement (13.4%), which greatly improves the ability of our model to find accurate correspondences and correct errors in geometry. We show qualitative results of our ablation in Figure 5. We evaluate the following model components: UF: Baseline 2D enhancer Uformer which our approach is based on. NN: Nearest Neighbour Selection. 3D-A: Our 3D Alignment module comprising of feature reprojection using depths and camera poses. NRF: Our Non-Rigid Refinement module comprising of an iterative lightweight optical flow network which further aligns the reprojected features. GA: Our Geometric Attention module comprising a geometry-aware feature regulation mechanism. PT: The effect of pre-training our model. We conduct a more detailed ablation study on pre-training and fine-tuning in Table 4. Here we show two things; that our method can achieve large improvements in as little as 1 minute of fine-tuning; and that it is our complete combined contributions which allow our method to learn novel image degradations so quickly. Uformer+NN also has access to nearest neighbours from the training set, but as it is unaware of geometry, it is not capable of quickly leveraging the information from nearby viewpoints.
@@
Limitations A limitation of our method is the requirement to store all relevant training images. This could be prohibitive for very large scenes, especially at higher image resolutions. Secondly, although we are several times faster than the baseline NeRF methods, our approach does not yet achieve real time (i.e. 30fps) inference. Societal Impact As NeRFs become more editable, our method could be misused to improve the photorealism of generated videos of people, places and objects. Currently our method requires ground truth images with a dense 3D coverage of the scene, so it is not trivial to use in the wild.
@@ 5 Conclusion
We have presented RoGUENeRF, a geometry-consistent NeRF enhancer which combines concepts from 3D and 2D vision to substantially improve the image quality of NeRF renderings in real world settings. Our model accurately finds correspondences between different camera views by performing 3D alignment and non-rigid refinement, while also being robust to errors in camera pose estimation and reducing reprojection artifacts with geometry-aware attention. RoGUENeRF achieves consistent improvements in image quality over six varied NeRF baselines and existing NeRF enhancers across three challenging real world datasets. Our model demonstrates wide applicability and strong generalization, fine-tuning on a novel scene in under minutes to learn the distribution of image degradations.